OpenStudy (anonymous):

y=(5x^2-1)^8(1-6x^3) can u help me using general formula

OpenStudy (anonymous):

general formula for what? you certainly do not want to multiply all this mess out!

OpenStudy (anonymous):

n.u^n-1 time DU/DX ... finding Y

OpenStudy (anonymous):

oooh you want the derivative right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

but can we use the derivative of quotient?

OpenStudy (anonymous):

\[y=(5x^2-1)^8(1-6x^3) \] use \[(fg)'=f'g+g'f\] with \[f(x)=(5x^2-1)^8,f'(x)=8(5x^2-1)^7\times 10x\\ g(x)=1-6x^3,g'(x)=-18x^2\]

OpenStudy (anonymous):

yesss .... then what next?

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

plug them in to the formula

OpenStudy (anonymous):

after pulging them ... what next?

OpenStudy (anonymous):

have a snack?

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] \[80x(5x^2-1)^7(1-6x^3)-18x^2(5x^2-1)^8\]

OpenStudy (anonymous):

i can't imagine trying to simplify this mess in any way, although both terms have a common factor of \((5x^2-1)^7\) which you could factor out if you like

OpenStudy (anonymous):

too much algebra for not enough payback i would leave it as it is